# Long Video Understanding with Learnable Retrieval in Video-Language Models

### v2

@@ Abstract
The remarkable natural language understanding, reasoning, and generation capabilities of large language models (LLMs) have made them attractive for application to video understanding, utilizing video tokens as contextual input. However, employing LLMs for long video understanding presents significant challenges. The extensive number of video tokens leads to considerable computational costs for LLMs while using aggregated tokens results in loss of vision details. Moreover, the presence of abundant question-irrelevant tokens introduces noise to the video reasoning process. To address these issues, we introduce a simple yet effective learnable retrieval-based video-language model (R-VLM) for efficient long video understanding. Specifically, given a question (query) and a long video, our model identifies and selects the most relevant $K$ video chunks and uses their associated visual tokens to serve as context for the LLM inference. This effectively reduces the number of video tokens, eliminates noise interference, and enhances system performance. We achieve this by incorporating a learnable lightweight MLP block to facilitate the efficient retrieval of question-relevant chunks, through the end-to-end training of our video-language model with a proposed soft matching loss. Our experimental results on multiple zero-shot video question answering datasets validate the effectiveness of our framework for comprehending long videos.
Index Terms:
Large Language Models, Long Video Question Answering, Retrieval-based Video-language Model.
@@ I Introduction
There is a growing trend towards leveraging large language models (LLMs) for video QA [1, 2, 3, 4]. On one hand, LLMs benefit from the vast knowledge acquired through training on enormous text corpora; on the other hand, they provide users with a more natural and intuitive way to interact with video data. Generally, visual tokens extracted from a video snippet are transformed and used as input (prompt) to the LLM, along with the text query. However, it is worth noting that the consumption of abundant visual tokens by a LLM can significantly increase the memory and computational burden, making it unaffordable for low-resource GPU agents. To mitigate this issue, Video-ChatGPT [1] performs global spatial and temporal pooling on the video tokens, although this comes at the cost of losing detail due to pooling. Video-LLaMA [4] aggregates video tokens using a Q-former with cross attention. Most of these methods are designed for short-video QA tasks, where answer-related frames usually spread over the trimmed video snippet.
@@ II Related Work
Some works have focused on enabling video-language understanding by integrating vision encoder and LLMs [23], [1], [3], [2]. Video-LLaMA [23] uses a video Q-former to assemble the features from the pre-trained CLIP image encoder. Video-ChatGPT [1] performs spatial and temporal pooling of the feature maps encoded by the pre-trained CLIP image encoder. The ensembled visual tokens are taken as input to the LLMs. Most of these works globally aggregate video tokens. They may work well for short videos. But for the long video-language comprehension, they would be less efficient. The information of the question-related video segments may be submerged to the global-wise token representations, making the reasoning difficult.
SeViLA [24] utilizes the BLIP-2 image-language model for selecting temporal key frames which are subsequently inputted into BLIP-2 for video-based question answering. The process involves complex training for convergence and the frame selection is computationally expensive due to the necessity of processing potential frames through the large BLIP-2 model (with even the smallest version being 3.1 billion parameters). This would not be suitable for long video understanding.
In order to understand long videos, MovieChat [25] leverages short-term and long-term memories for video representation. LLaMA-VID [26] represents every frame coarsely by two tokens to reduce the computation and memory cost for long video reasoning, which however still introduces temporal redundancy and interference in LLM inference. SeViT [27] proposes a framework for long untrimmed video understanding. It includes a non-parametric frame retriever to select a few query-relevant frames from the video and uses them as the input to a video-grounded text generator. However, the frame selector is non-parametric and the system cannot be end-to-end optimized, resulting in the inferiority of the selected frames and sub-optimal performance. Different from SeViT, we introduce learnable retrieval through a light-weighted MLP block for building an end-to-end video-language model, powered by LLM. In [28], a Frame-Prompter is designed to select key video frames. However, with only video frames as input to the Frame-Prompter, the selection is not question aware and would lead to ineffective. Wang et al.[29] fuse the question-and-answer pairs as event descriptions to find keyframes as target moments and pseudo-labels. Unlike [29], we do not use pseudo-labels for supervision. Our framework leverages the end-to-end learning to automatically select the question-relevant chunks, providing a large optimization space and avoiding the side effects from the inaccuracy of pseudo-labels. VideoTree [30] extracts query-relevant information from the input video through an iterative process, based on their relevance to the query scored by an LLM. However, leveraging an additional LLM for key frame selection increases the complexity of the system. In contrast, our retrieval design is lightweighted and end-to-end trained in the VLM.
@@ III-A Video Tokenization
[…]
@@
Chunk wise compression The preliminary token number of a chunk is large, which would result in a high computational burden and large memory requirement for the LLM even though we only select a few chunks as input to LLM. Therefore, we perform chunk-wise pooling to obtain reduced spatial and temporal tokens (see Fig. 3 for the illustration in Supplementary Section A).
@@ III-B Question-guided Retrieval for Chunk Selection
where the MLP block is a two-layer perceptron consisting of two fully connected layers and an ReLU activation function in between. This MLP transformation strengthens the correlation between the question representation and the potential corresponding chunk features for chunk selection. The MLP block is very lightweight with negligible number of parameters (i.e., 1.8M), while facilitating the question-aware retrieval for long video understanding. Note that achieving this is non-trivial, as we do not have any ground-truth locations of the question relevant chunks for supervision. The purpose of adding an MLP block to the text features is to enable learnable retrieval, i.e., the learnable selection of question-related video chunks. Another alternative design is to apply an MLP to the vision features during retrieval or apply MLPs to both. For a video question pair in inference, if we apply the MLP to the text encoder, we only need the inference through the MLP once. If we apply the MLP to the vision encoder, we need the inference through the MLP for every chunk. For simplicity, we only apply to the text feature.
@@ III-C End-to-End Optimization
where $\bar{\mathbf{v}}_{i}$ is a weighted combination of the $\mathcal{L}_{i}$ chunk features. For a weight $e^{s_{i}^{j}}$, $e$ denotes the base of the natural logarithm, $s_{i}^{j}$ denotes the similarity score between the query and the $j^{th}$ chunk features (see Eq. (2)) of the video $V_{i}$. $\mathbf{v}_{i}^{j}$ denotes the averaged token feature of the $j^{th}$ chunk, i.e., $\mathbf{v}_{i}^{j}=\frac{1}{N+M}\sum_{n=1}^{N+M}\mathbf{u}_{i,n}^{j}$. The soft matching loss aims to align the aggregated visual feature with the question feature by adjusting weights. Through the optimization of the MLP block, high weights are assigned to video chunks closely related to the question and low weights to less relevant chunks, thus improving the selection of relevant chunks for accurate question answering.
@@
where $\mathcal{L}^{pred}_{i}$ denotes the cross-entropy loss between the generated prediction and groundtruth answer (a.k.a. LLM loss). $\lambda$ is a hyper-parameter (we set to 10) that balances the contribution of the regularization term (see Table X for the ablation).
@@ IV-A Implementation Details
Following Video-ChatGPT [1], we use LLaVA [17] as our base model. We utilize the pre-trained CLIP ViT-L/14 [20] as our image encoder and extract the feature from the second-to-last layer as the $h\times w$ visual tokens of a frame. We use the fine-tuned Vicuna (7B) from LLaVA as our LLM. For the text encoder, we use the pre-trained CLIP ViT-L/14 text encoder and extract the class token feature of the penultimate layer. Each of the two fully connected layers within the MLP block is equipped with 1024 neurons. The projector has 4096 neurons.
@@ IV-B Datasets and Evaluation
TABLE I: Information of the evaluation datasets, including the number of videos, the number of QA pairs, and the average video duration.
@@
We use Video Instruction Data collected by Maaz et al.[1] for video instruction tuning. This dataset contains about 100k question and answer pairs based on the Activitynet-QA dataset (average duration 180 seconds), including various types of questions.
We evaluate the generalization (zero shot) performance of our framework on long video QA datasets Activitynet-QA [31], EgoSchema [32], NExt-QA[33], and IntentQA [34]. To widely evaluate the effectiveness of our model, besides the four popular datasets, we also tested on other long video QA datasets WildQA [35], QaEgo4D [36], lifeQA [37], and Social-IQ 2.0 [38]. Table I shows specific information about each dataset. We perform ablation studies on Activitynet-QA, WildQA, QaEgo4D, lifeQA, and Social-IQ 2.0. Note that the other widely used datasets MSVD-QA [39], MSRVTT-QA [40] for video QA are not suitable, which have average duration of only 10 seconds and 15 seconds, being too short for evaluation.
We assess model performance using accuracy and a 0-5 similarity score, respectively, as per Video-ChatGPT [1], with ChatGPT (gpt-3.5-turbo) evaluating the correctness of predictions through binary and numerical judgments. See Supplementary Section B-A for more information.
@@
TABLE VI: Comparison with the other video-language models, including Video-LLaMA [23] and the baseline method Video-ChatGPT [1]. We report the accuracy (%)/ average score.
@@ IV-C Comparison with Other Models
Table II, III, IV, and Table VI show the performance comparison on Activitynet-QA, EgoSchema, NExt-QA, and IntentQA, respectively. We can see that our model achieves the state-of-the-art performance. On Activitynet-QA, our R-VLM outperforms LLaMA-VID-7B by 2.3% in accuracy. Our R-VLM achieves competitive performance with VideoChat2, even though VideoChat2 uses much more training data than ours (1.1 million vs. our 100K video-text pairs).
On the WildQA, QaEgo4D, lifeQA, and Social-IQ 2.0 datasets, Table VI shows the comparison with the baseline method Video-ChatGPT, and the method of Video-LLaMA [23]. R-VLM outperforms Video-ChatGPT significantly by 6.8%, 2.8%, 4.8%, 6.0% in accuracy, respectively. Ours consistently achieves the best average score. Note that the number of visual tokens of our R-VLM is comparable to that of Video-ChatGPT (i.e., 340 vs. 356), making a fair comparison. This demonstrates the effectiveness of our learnable retrieval-based design, which facilitates the exploration of the most informative visual tokens and preservation of necessary vision details for QA. We found that when evaluating Video-LLaMA, the metric of accuracy is not as reliable as the score. Video-LLaMA is prone to give detailed description of the entire video, whereas the answer is usually not question-specific. The accuracy metric treats such answer as correct, while the score rating is relatively reasonable. Fig. 5 in Supplementary Material B-E shows a typical example on QaEgo4D, where the answer includes much irrelevant (or incorrect) information while the answer is submerged in the overall answer. In contrast, our model provides more concise and accurate answers. This explains why our accuracy is lower than Video-LLaMA on QaEgo4D.
footnotetext: Sevilla [24] achieves an average accuracy of 73.8%, where the model is trained on NExT-QA. In contrast, our model has not seen Next-QA during training. VideoChat2 [48] achieves an average accuracy of 68.6%. It uses much more data for training (1.9M samples vs. ours using 100K).
TABLE VII: Comparison of different methods and ablation studies. All these models are trained using the same video instruction data. R-VLM denotes our final scheme with learnable retrieval. Video-ChatGPT is our baseline method. R-VLM w/ Uni. denotes uniform sampling of $K$ chunks in our framework instead of retrieval-based sampling. R-VLM w/ CLIP M. denotes that we use the final CLIP class token features of vision and text for matching in our framework, without incoperating learnable parameters for retrieval. R-VLM w/o SM denotes that we do not use the proposed soft matching (SM) loss. R-VLM w/o G. denotes that the gradient back-propagation from the LLM loss $\mathcal{L}^{pred}_{i}$ to the learnable MLP block $\psi$ is disabled. We report the accuracy (%) / average score.
@@
IV-D Ablation Studies
@@
In this section, we study and validate the effectiveness of our retrieval designs, the soft matching loss, chunk-wise design, the influence of hyperparameter $K$, and the influence of the time duration of a chunk, respectively.
Effectiveness of Retrieval for Chunk Selection Within our framework, we compare our retrieval mechanism and the uniform sampling strategy for the selection of $K=5$ chunks as context input to the LLM. For the uniform sampling setting which we name as R-VLM w/ Uni., the model selects $K$ from $N$ video chunks uniformly instead of based on question guided retrieval. Table  VII shows that our final scheme R-VLM with learnable retrieval-based strategy outperforms R-VLM w/ Uni. by 3.6%, 0.9%, 2.2%, and 5.7% on WildQA, QaEgo4D, lifeQA, and Social-IQ 2.0, respectively. For a long video, the video chunks relevant to the question usually account for a small portion of the entire video. It is difficult to hit these question-related chunks with uniform sampling. A large language model may not correctly answer a question when it accepts video chunks not correlated with the question. In contrast, our learnable retrieval learns to select the chunks most relevant to the question, providing more reliable and less redundant information to the LLM for effective inference. Visualizations of the selected chunks in Fig. 2 also validate this and show that our retrieval-based model is interpretable, indicating from where and why the model produces the answer.
@@
Effectiveness of Learnable Retrieval vs. Off-the-shelf CLIP based Retrieval We propose the learnable retrieval by introducing the MLP block. One may wonder the performance of using off-the-shelf CLIP class token features to match question and chunk features inside our framework. To answer this, we design a scheme R-VLM w/ CLIP M. for comparison. For a chunk, we use the averaged class token feature of the last layer of the CLIP image encoder as the vision chunk feature, and the class token feature of the last layer of the CLIP text encoder as the question feature for matching. There is no learnable parameters for the retrieval. Table VII shows that our R-VLM with learnable retrieval consistently outperforms R-VLM w/ CLIP M., upto 7.3% in accuracy on the lifeQA dataset. That may be because the CLIP matching is originally designed for image and caption matching, which is not robust to the image and question matching. Thanks to the adaptation of the text encoder through a learnable MLP block, our question guided retrieval can better identify the relevant video chunks.
@@
Influence of Soft Matching (SM) Loss In order to regularize the learning of the MLP block for better retrieval, we introduce SM loss. Table VII shows the comparison of our framework without the SM loss (R-VLM w/o SM) and that with SM loss (R-VLM). We can see that incorporating the SM loss significantly improves the performance. Our final scheme with the SM loss i.e., R-VLM outperforms that without SM Loss. SM Loss uses the cosine similarity between text embedding and video tokens as weight to re-weight the video tokens to obtain new video tokens. Then maximize the similarity between the new video tokens and the text embedding. This facilitates our learnable retrieval layer to find the video clips most similar to the questions.
@@
Influence of LLM Loss to Retrieval The instruction tuning with the LLM loss ($\mathcal{L}^{pred}_{i}$) back-propagates the gradient also to the learnable MLP block $\psi$, which facilitates the selection of right chunks for retrieval. When we disable the back-propagation to $\psi$, we have scheme R-VLM w/o G.. As shown in Table VII, our final scheme R-VLM that allows the back-propagation obviously outperforms R-VLM w/o G. by 2.6%, 0.9%, 1.6%, and 1.2% on WildQA, QaEgo4D, lifeQA, and Social-IQ 2.0.
@@
Influence of Hyperparameter $K$ In general, when $K$ is too small, it may lead to a loss of information necessary to answer the question. When $K$ is too large, interference may be introduced, confusing the LLM. Table VIII shows the results of using different $K$. We found that as $K$ gradually increases from 1 to 5, the average performance increases. When $K$ increases from 5 to 7, the performance decreases. We found $K=5$ presents a good trade-off on most datasets, even though there are slight differences on different datasets. We leave the adaptive design of $K$ as future work.
TABLE VIII: Ablation study on the influence of $K$, evaluated in terms of accuracy (%)/score. We use bold to mark the best performance and underline to mark the second-best performance.

### v3

@@ Abstract
The remarkable natural language understanding, reasoning, and generation capabilities of large language models (LLMs) have made them attractive for application to video understanding, utilizing video tokens as contextual input. However, employing LLMs for long video understanding presents significant challenges. The extensive number of video tokens leads to considerable computational costs for LLMs while using aggregated tokens results in loss of vision details. Moreover, the presence of abundant question-irrelevant tokens introduces noise to the video reasoning process. To address these issues, we introduce a simple yet effective learnable retrieval-based video-language model (R-VLM) for efficient long video understanding. Specifically, given a question and a long video, our model identifies the most relevant $K$ video chunks and uses their associated visual tokens to serve as context for the LLM inference. This effectively reduces the number of video tokens, eliminates noise interference, and enhances system performance. We achieve this by incorporating a learnable lightweight MLP block to facilitate the efficient retrieval of question-relevant chunks, through the end-to-end training of our video-language model with a proposed soft matching loss. Experimental results on multiple zero-shot video question answering datasets validate the effectiveness of our framework for comprehending long videos.
@@ I Introduction
There is a growing trend towards leveraging LLMs for video QA [1, 2, 3, 4, 5]. On one hand, LLMs benefit from the vast knowledge acquired through training on enormous text corpora; on the other hand, they provide users with a more natural and intuitive way to interact with video data. Generally, visual tokens extracted from a video snippet are transformed and used as input (prompt) to the LLM, along with the text query. However, it is worth noting that the consumption of abundant visual tokens by a LLM can significantly increase the memory and computational burden, making it unaffordable for low-resource GPU agents. To mitigate this issue, Video-ChatGPT [1] performs global spatial and temporal pooling on the video tokens, although this comes at the cost of losing detail due to pooling. Video-LLaMA [4] aggregates video tokens using a Q-former with cross attention. Most of these methods are designed for short-video QA tasks, where answer-related frames usually spread over the trimmed video snippet.
@@ II Related Work
Some works have focused on enabling video-language understanding by integrating vision encoder and LLMs [24, 1, 3, 2],[25]. Video-LLaMA [24] uses a video Q-former to assemble the features from the pre-trained CLIP image encoder. Video-ChatGPT [1] performs spatial and temporal pooling of the feature maps encoded by the pre-trained CLIP image encoder. The ensembled visual tokens are taken as input to the LLMs. Most of these works globally aggregate video tokens. They may work well for short videos. But for the long video-language comprehension, they would be less efficient. The information of the question-related video segments may be submerged to the global-wise token representations, making the reasoning difficult.
SeViLA [26] uses the BLIP-2 image-language model to select temporal key frames which are subsequently inputted into BLIP-2 for video question answering. The process involves complex training and the frame selection is computationally expensive since the candidate frames must be processed by BLIP-2 (minimum 3.1B parameters), making it inefficient for long video understanding.
In order to understand long videos, MovieChat [27] leverages short- and long-term memories for video representation. LLaMA-VID [28] reduces the computation and memory cost by representing each frame with two tokens, but still suffers from temporal redundancy and interference. EALD-MLLM [29] addresses long video emotion analysis by segmenting videos into sampled clips, generating a description for each clip with Video-LLaMA, and leveraging ChatGPT to aggregate segment-level outputs into coherent global predictions. VILA [30] employs a Frame-Prompter for key-frame selection, but without question awareness, the selection is ineffective. SeViT [31] retrieves query-relevant frames for a text generator, yet its non-parametric frame selector prevents end-to-end optimization, leading to sub-optimal performance. Wang et al.[32] fuse QA pairs as event descriptions to find key-frames as target moments and pseudo-labels. Unlike [32], we do not use pseudo-labels for supervision. VideoStreaming [33] splits video into short clips, where a small language model (Phi-2, 2.7B) was used for clip selection. VideoTree [34] extracts query-relevant information from the input video through an iterative process, based on their relevance to the query scored by an LLM. However, requiring an extra LLM adds complexity. In contrast, our method introduces a lightweight end-to-end trainable retrieval module with Soft Matching loss design for efficient video understanding. Such lightweight, efficient, and easy to integrate design is suitable for practical system deployment.
@@ III-A Video Tokenization
Chunking the video and feature extraction Given a long video sample $V_{i}\in\mathbb{R}^{T_{i}\times H\times W\times 3}$ with $T_{i}$ frames, height $H$, and width $W$, we divide it into small non-overlapped video chunks (see Fig. 1), which are the basic units with each containing spatial information and temporal dynamic for question-guided retrieval. We set the duration of a chunk to 4 seconds, i.e., $M=4$ frames when frame rate is 1 frame per second (fps). We have $L_{i}=\lceil T_{i}/M\rceil$ chunks, where $\lceil\cdot\rceil$ denotes the ceiling function.
@@
Chunk wise compression The preliminary token number of a chunk is large, which would result in a high computational burden and large memory requirement for the LLM even though we only select a few chunks as input to LLM. Therefore, we perform chunk-wise pooling to obtain reduced spatial and temporal tokens (see Fig. 2 for the illustration in Supplementary Section VI).
@@ III-B Question-guided Retrieval for Chunk Selection
Note that achieving this is non-trivial, as we do not have any ground-truth locations of the question relevant chunks for supervision. We propose learnable retrieval, i.e., the learnable selection of question-related video chunks, through end-to-end training. Particular, we add an MLP block to the text features to facilitate the matching. Another alternative design is to apply an MLP to the vision features during retrieval or apply MLPs to both the text encoder and the visual encoder. For a video question pair in inference, if we apply the MLP to the text encoder, we only need the inference through the MLP once. If we apply the MLP to the vision encoder, we need the inference through the MLP for every chunk. For simplicity, we apply to the text feature.
@@
where the MLP block is a two-layer perceptron consisting of two fully connected layers and an ReLU activation function in between. This MLP transformation strengthens the correlation between the question representation and the potential corresponding chunk features for chunk selection. The MLP block is very lightweight with negligible number of parameters (i.e., 1.8M), while facilitating the question-aware retrieval for long video understanding.
@@ III-C End-to-End Optimization
where $\bar{\mathbf{v}}_{i}$ is a weighted combination of the $\mathcal{L}_{i}$ chunk features. For a weight $^{s_{i}^{j}}, ee$ denotes the base of the natural logarithm, $s_{i}^{j}$ denotes the similarity score between the query and the $j^{th}$ chunk features (see Eq. (2)) of the video $V_{i}$. $\mathbf{v}_{i}^{j}$ denotes the averaged token feature of the $j^{th}$ chunk, i.e., $\mathbf{v}_{i}^{j}=\frac{1}{N+M}\sum_{n=1}^{N+M}\mathbf{u}_{i,n}^{j}$. The soft matching loss aims to align the aggregated visual feature with the question feature by adjusting weights. Through the optimization of the MLP block, high weights are assigned to video chunks closely related to the question and low weights to less relevant chunks, thus improving the selection of relevant chunks for accurate question answering.
@@
where $\mathcal{L}^{pred}_{i}$ denotes the cross-entropy loss between the generated prediction and groundtruth answer (a.k.a. LLM loss). $\lambda$ is a hyper-parameter (we set to 10) that balances the contribution of the regularization term (see Table XIII for the ablation).
@@ IV-A Implementation Details
Following Video-ChatGPT [1], we use LLaVA-7B [18] with CLIP ViT-L/14 [21] as both image and text encoders. The MLP block has two FC layers of 1024 neurons, and the projector has 4096 neurons. We fine-tune only the MLP and projector, keeping encoders and LLM frozen. With $K=5$, we obtain 340 visual tokens, comparable to Video-ChatGPT’s 356. Unlike Video-ChatGPT, which trains only a 4.2M-parameter projector, we also train a 1.8M-parameter MLP, totaling 6.0M trainable parameters. Please see Supplementary Section VII-A for more details.
@@ IV-B Datasets and Evaluation
TABLE I: Information of the evaluation datasets, including the number of videos, QA pairs, and the average video duration.
@@
We use Video Instruction Data collected by Maaz et al.[1] for video instruction tuning. This dataset contains about 100k question and answer pairs based on the Activitynet-QA dataset (average duration 180 seconds).
We evaluate the generalization (zero shot) performance of our framework on long video QA datasets Activitynet-QA [35], EgoSchema [36], NExt-QA[37], and IntentQA [38]. To widely evaluate the effectiveness of our model, besides the four popular datasets, we also tested on other long video QA datasets WildQA [39], QaEgo4D [40], lifeQA [41], and Social-IQ 2.0 [42]. Table I shows specific information about each dataset. We perform ablation studies on Activitynet-QA, WildQA, QaEgo4D, lifeQA, and Social-IQ 2.0.
We evaluate performance using accuracy and a 0–5 similarity score following Video-ChatGPT [1] (see Supplementary Section VII-B for details).
@@
TABLE VI: Comparison with the other video-language models, including Video-LLaMA [24] and the baseline method Video-ChatGPT [1] (accuracy (%)/ average score).
@@ IV-C Comparison with Other Models
Table II, III, IV, and Table V show the performance comparison on Activitynet-QA, EgoSchema, NExt-QA, and IntentQA, respectively. Note that EgoSchema, NExt-QA, and IntentQA were not involved in training but are only used for testing. Our R-VLM achieves the state-of-the-art performance. On Activitynet-QA, our R-VLM outperforms LLaMA-VID-7B by 2.3% in accuracy. Our R-VLM achieves competitive performance with VideoChat2, even though VideoChat2 uses much more training data than ours (1.1 million vs. our 100K video-text pairs).
On WildQA, QaEgo4D, lifeQA, and Social-IQ 2.0, Table VI shows the comparison with the baseline method Video-ChatGPT, and the method of Video-LLaMA [24]. R-VLM outperforms Video-ChatGPT significantly by 6.8%, 2.8%, 4.8%, 6.0% in accuracy, respectively. Ours consistently achieves the best average score. Note that the number of visual tokens of our R-VLM is comparable to that of Video-ChatGPT (i.e., 340 vs. 356), making a fair comparison. This demonstrates the effectiveness of our learnable retrieval-based design, which facilitates the exploration of the most informative visual tokens and preservation of necessary vision details for QA. We found that when evaluating Video-LLaMA, the metric of accuracy is not as reliable as the score. Video-LLaMA is prone to give detailed description of the entire video, whereas the answer is usually not question-specific. The accuracy metric treats such answer as correct, while the score rating is relatively reasonable. Fig. 4 in Supplementary Section VII-F shows a typical example on QaEgo4D, where the answer includes much irrelevant (or incorrect) information while the answer is submerged in the overall answer. In contrast, our model provides more concise and accurate answers. This explains why our accuracy is lower than Video-LLaMA on QaEgo4D.
TABLE VII: Comparison (accuracy (%) / average score) of different methods and ablation studies. All these models are trained using the same video instruction data. R-VLM denotes our final scheme with learnable retrieval. Video-ChatGPT is our baseline method. R-VLM w/ Uni. denotes uniform sampling of $K$ chunks in our framework instead of retrieval-based sampling. R-VLM w/ CLIP M. denotes that we use the final CLIP class token features of vision and text for matching without incorporating learnable parameters for retrieval. R-VLM w/o SM denotes that we do not use the proposed soft matching (SM) loss. R-VLM w/o G. denotes that the gradient back-propagation from the LLM loss $\mathcal{L}^{pred}_{i}$ to the learnable MLP block $\psi$ is disabled.
@@
Average 45.43/2.96 47.33/2.98 46.66/2.94 46.59/2.92 48.35/3.01 49.87/3.03
IV-D Complexity Analysis
Section VII-D in Supplementary provides detailed analysis.
IV-E Ablation Studies
@@
We study the effectiveness of our retrieval designs, the soft matching loss, chunk-wise design, the influence of hyperparameter $K$, comparison with Q-Former-based designs, etc.
Effectiveness of Retrieval for Chunk Selection We compare our retrieval mechanism and the uniform sampling strategy for selecting $K=5$ chunks as LLM input. In the uniform sampling setting (R-VLM w/ Uni.), we select $K$ chunks uniformly rather than by question guided retrieval. Table  VII shows our scheme R-VLM with learnable retrieval outperforms R-VLM w/ Uni. by 3.6%, 0.9%, 2.2%, and 5.7% on WildQA, QaEgo4D, lifeQA, and Social-IQ 2.0, respectively.
Since question-relevant chunks usually form only a small portion of a long video, uniform sampling often misses them, leading the LLM to wrong answers. Our learnable retrieval instead selects the most relevant chunks, providing reliable, non-redundant context. Visualizations in Fig. 4 further show its interpretability by indicating which chunks support the answer.
@@
Effectiveness of Learnable Retrieval vs. Off-the-shelf CLIP based Retrieval We introduce learnable retrieval via the MLP block. To compare, we design R-VLM w/ CLIP M. that which uses the averaged class token feature of the last layer of the CLIP image encoder as the vision chunk feature, and the class token feature of the last layer of the CLIP text encoder as the question feature for matching, with no learnable parameters. Table VII shows our R-VLM with learnable retrieval consistently outperforms R-VLM w/ CLIP M., up to 7.3% on LifeQA. This is likely because CLIP is trained for image-caption matching, whereas our MLP-adapted retrieval better aligns question and relevant video chunks.
@@
Influence of Soft Matching (SM) Loss SM loss regularizes MLP learning for better retrieval. Table VII shows our framework with SM loss (R-VLM) outperforms the version without it (R-VLM w/o SM). SM loss re-weights video chunk tokens based on cosine similarity with the text embedding, then maximizes similarity between the video chunk tokens and the text embedding. This facilitates our learnable retrieval layer to identify the video clips most related to the questions.
@@
Influence of LLM Loss to Retrieval The instruction tuning with the LLM loss ($\mathcal{L}^{pred}_{i}$) also propagates the gradient to the learnable MLP block $\psi$, which facilitates the selection of right chunks for retrieval. When we disable the back-propagation to $\psi$, we have scheme R-VLM w/o G.. Table VII shows our final scheme R-VLM that allows the back-propagation obviously outperforms R-VLM w/o G. by 2.6%, 0.9%, 1.6%, and 1.2% on WildQA, QaEgo4D, lifeQA, and Social-IQ 2.0.
@@
Influence of Hyperparameter $K$ A small $K$ may miss necessary information, while a large $K$ may introduce interference. Table VIII shows that as $K$ gradually increases from 1 to 5, the average performance increases. When $K$ increases from 5 to 7, the performance decreases. We found $K=5$ presents a good trade-off on most datasets, even though there are slight differences on different datasets. In addition, we calculated the computation (FLOPs) cost saving (measured by percentage) using different $K$, compared with that of using all chunks (see Section VII-D in Supplementary). Table VIII shows that given $K=5$, the percentage of saving ranges from 64% to 95% for different datasets (due to different video lengths), demonstrating the high efficiency of our solution. For ActivityNet-QA, the percentage of saving ranges from 93% to 73% when $K$ is set to 1 to 7. We leave the adaptive design of $K$ as future work.
TABLE VIII: Ablation study on the influence of $K$, evaluated in terms of accuracy (%) /score /computation-saving (%) in LLM inference. We use bold to mark the best performance and underline to mark the second-best performance.